# Sharp transitions in small exciton spectra for multi-orbital lattice systems

### v2

@@ Abstract
We demonstrate that strongly bound excitons, whose radii approach the lattice constant, display physics that eludes continuum descriptions not just quantitatively but qualitatively. We investigate such phenomenology by calculating the exciton spectra for several one- and two-dimensional lattice models, whose valence band has a multi-orbital character. We identify sharp transitions in the character and momentum of the lowest-energy exciton, driven directly by the multi-orbital nature of the lattice models. Such transitions cannot occur in simple continuum descriptions.
@@ I Introduction
Excitons—electron-hole pairs bound by their mutual Coulomb attraction—are observed in the optical response of most semiconductors. In conventional inorganic semiconductors such as Si, the strength of the attraction is small compared to the bandwidths of the conduction and valence bands, and the resulting Wannier-Mott excitons [1, 2, 3, 4] are weakly bound and have a radius that greatly exceeds the lattice constant $a$. Consequently, a continuum description suffices: the maximum and minimum of valence and conduction bands are well approximated as quadratic, describing well the exciton wavefunction which is dominated by the contribution of the optima of these bands. For three-dimensional (3D) crystals and a $1/r$ Coulomb attraction, this maps the exciton spectrum to that of a hydrogen atom, with a rescaled Bohr radius and Rydberg energy [2, 3, 4, 5, 6].
@@
For a screened, short-range attraction with a characteristic scale $U$, excitons appear in 3D only if $U$ exceeds a critical value. However, both in two dimensions (2D) and one dimension (1D), a bound state exists for arbitrarily small $U$, with a binding energy $E_{\text{bind}}\propto\exp(-W/U)$ and $E_{\text{bind}}\propto U^{2}/W$, respectively [5, 8, 7] (here $W$ is a measure of the average kinetic energy of the pair). These asymptotic estimates are widely used in standard many-body approaches to excitons [9, 10] and provide a baseline for comparison between theory and experiments, for example, see Refs. 11, 12.
@@
However, many low-dimensional semiconductors that are technologically relevant, host strongly bound Frenkel and charge-transfer excitons [13, 14, 15] whose radii are equal or comparable to the lattice constant $a$. For example, TR-ARPES on epitaxial C60 films on Au(111) has revealed binding energies of 300–1000 meV for excitons that are either on the same C60 molecule, or have the electron and hole on neighbor molecules [17, 16, 18, 19, 20]. These results agree with earlier work on C60 using other optical techniques [21, 22]. Other examples include organic crystals such as pentacene and PTCDA, which also exhibit exciton binding energies on the order of 0.5 eV [23]. In monolayer WS2, two-photon spectroscopy has measured an exciton binding energy of $\sim 0.3$ eV, underlining that atomically thin semiconductors host non-hydrogenic excitons [24]. For these small excitons, the specific lattice symmetry becomes relevant beyond just the renormalization of the conduction and valence band effective masses [25]. This is because the strong attraction mixes within the exciton wavefunction hole/electron states lying far below/above the top/bottom of the valence/conduction band, and their dispersion is not described well by the continuum approximation. Furthermore, especially in organic semiconductors, the conduction and valence bands often have multi-orbital character, leading to distinct spectral fingerprints and momentum-dependent symmetry changes [26, 27, 28] which may be completely missed by a continuum approximation.
These considerations show that the convenient, simple estimates provided by the continuum approximation must become quantitatively inaccurate once small excitons are concerned, yet it is generally believed that they remain qualitatively correct. In this work, we show that it is very easy to find examples where these simple estimates are qualitatively wrong, especially for small excitons in multi-orbital models. Fortunately, solving such two-particle problems (here, the electron and hole) on an infinite lattice can be done very efficiently, as we demonstrate with a method adapted from studies of few-particle interacting systems, see e.g. Ref. 29. We hope that such methods will be adopted to study realistic lattice models and obtain accurate properties of their small excitons.
@@
The article is organized as follows: Section II introduces the lattice models we study, and Section III details our formalism. Our results are shown in Section IV, and we summarize our findings in Section V.
@@ II The lattice models
Here $n,m$ index lattice sites located at ${\bf R}_{n},{\bf R_{m}}$. $\alpha,\beta$ label the specific valence orbitals, and $v^{\dagger}_{\alpha n}$ creates an electron in the orbital $\alpha$ at site $n$. Explicitly, we will show results for 1D models with 2 orbitals per site ($1s$ and $2p_{x}$ for a chain parallel to the $x$ axis) and 2D triangular models with 3 orbitals per site ($d_{x^{2}-y^{2}},d_{xy},d_{3z^{2}-r^{2}}$ for a lattice lying in the $xy$ plane). Any other combinations of orbitals can be studied similarly. Crystal fields can be included trivially, but we ignore them for simplicity. We restrict the hopping to nearest-neighbor for simplicity, but this can be generalized straightforwardly to longer range. To minimize the number of free parameters for the 2D system, we use the Harrison rules [30] to calculate ratios between the hopping integrals, see Appendix A for details. For the same reasons, we present results only for a single-orbital conduction band:
@@
where $c_{n}^{\dagger}$ creates a conduction-band electron at site $n$ of the lattice. Here $\Delta$ is the difference between the on-site energies, and controls the value of the band gap. For simplicity we consider again only nearest-neighbor hopping; generalization to longer range hopping is straightforward. Similarly, generalizations to a multi-orbital conduction band is straightforward but introduces more free parameters. We note that in this work we ignore possible hybridization between the orbitals defining the conduction and valence bands. These can be included but lead to more significant complications, in particular lifting of the degeneracy between singlet and triplet excitons. In their absence, the spin degrees of freedom are trivial, thus we drop the spin labels in this work.
@@
At first sight, this on-site attraction is an unreasonable approximation for organic semiconductors whose small excitons are due precisely to the very weakly screened, long-range Coulomb interaction. Nevertheless, in the limit of on-site Frenkel excitons of most interest to us in this work, the electron and hole are on the same site with extremely high probability. Because they never 'explore' longer separations, the details of the longer-range Coulomb attraction become irrelevant. The only meaningful interaction energy scale in this limit is the difference between the attraction between an on-site pair and one occupying adjacent sites. This is what the $U_{\alpha}$ energies describe. For completeness, we mention that longer range Coulomb attraction can also be studied with the method we discuss below. A sample calculation for the effect of the nearest neighbor interaction can be found in Appendix C.
@@ III The method
Because of the invariance to translations of the Hamiltonian (1), the exciton momentum $\bf K$} is a good quantum number; its value is restricted to the Brillouin zone of the appropriate lattice. For any value of $\bf K$}, a particularly convenient basis for calculating exciton eigenstates is the set of all states [2, 9, 29]:
$|{\alpha,{\bf K},\bm{\delta}\rangle=\sum_{n}e^{i{\bf K}({\bf R}_{n}+\bm{\delta}/2)}c^{\dagger}_{n}v_{\alpha,n+\bm{\delta}}|{\rm GS}\rangle}$ (7)
where $|{\rm GS}\rangle$ is the ground state with filled valence bands and empty conduction band. Instead of using this basis to solve the Schrödinger equation, it is more convenient to calculate instead the particle-hole propagators [29]:
@@
$G_{\alpha,\beta}({\bf K},\bm{\delta},z)=\sum_{i}\frac{\langle\alpha,{\bf K},\bm{0}|{\bf K},i\rangle\langle{\bf K},i|\beta,{\bf K},\bm{\delta}\rangle}{z-E({\bf K},i)}$ (9)
@@
In the absence of Coulomb interactions, these electron-hole eigenstates are $c^{\dagger}_{\bf K+k}v_{{\bf k},\gamma}|{\rm GS}\rangle$ for any $\bf k$}, with the corresponding energy $E_{c}({\bf K+k})-E_{v}({\bf k},\gamma)$. They define the electron-hole continuum that for a given $\bf K$} starts at
$E_{\rm gap}({\bf K})=\min_{{\bf k},\gamma}[E_{c}({\bf K+k})-E_{v}({\bf k},\gamma)]$ (10)
@@
If the Coulomb attraction is sufficiently strong to bind excitons (which is always the case in 1D and 2D systems), their eigenenergies are signalled by discrete poles of $G_{\alpha,\beta}({\bf K},\bm{\delta},z)$ at energies inside the gap $\omega=E_{\rm exc}({\bf K})<E_{\rm gap}({\bf K})$. The corresponding exciton binding energy is $E_{\rm bind}({\bf K})=E_{\rm gap}({\bf K})-E_{\rm exc}({\bf K})$ [4, 32].
@@
meaning that
$\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle\propto G_{\alpha,\beta}({\bf K},\bm{\delta},E_{\rm exc}({\bf K})).$ (12)
Eq. (12) gives the real-space exciton amplitude, which we visualize after normalization with the condition $\sum_{\beta,\bm{\delta}}|\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle|^{2}=1$ (Fig. 7,8). We also use it to extract a characteristic exciton radius $\xi$ from the long-range envelope of the eigenstate $\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle\propto\exp(-|\bm{\delta}|/\xi)$ when $|\bm{\delta}|\rightarrow\infty$ [33].
These propagators are calculated from matrix elements between states of Eq. (7) of the identity $\hat{G}(z)(z-{\cal H})=1$, from which we find:
@@
The second term can be expanded in terms of other propagators, leading to a linear system of coupled equations involving all these propagators. Although in principle the size of this system is infinite because $\bm{\delta}$ can be arbitrarily large, in practice the magnitude of the propagators decreases very fast [33] with increasing $\bm{\delta}$ for in-gap energies comparable to those of small excitons. This allows us to truncate this system by setting to zero all propagators with $|\bm{\delta}|$ larger than a cutoff, which is chosen such that the exciton energy converges within the desired accuracy. For very small excitons, good accuracy can be obtained even for a cutoff set to $2a$ (see example below) making such calculations extremely efficient. Full details and equations are given in Appendix B.

### v3

# Sharp transitions in the spectra of small Frenkel-like excitons for multi-orbital lattice systems

@@ Abstract
We propose a method for calculating exciton spectra and wavefunctions for model lattice Hamiltonians, based on real-space electron-hole propagators. We verify that our results agree with those of the continuum approximation in the limit of large Wannier excitons, and propose a simple criterion to estimate the exciton size above which the continuum approximation is quantitatively accurate. We then investigate simple one- and two-dimensional multi-orbital lattice models and show that small, Frenkel-like excitons, whose size approaches the lattice constant, can display physics that disagrees with the simplest continuum descriptions (a single-valley quadratic expansion around the minimum gap) not just quantitatively, but qualitatively. Specifically, we identify sharp transitions in the character and momentum of the lowest-energy exciton, enabled by the multi-orbital nature of the lattice models.
@@ I Introduction
Excitons – electron-hole pairs bound by their mutual Coulomb attraction—are observed in the optical response of most semiconductors. In conventional inorganic semiconductors such as Si, the strength of the attraction is small compared to the bandwidths of the conduction and valence bands, and the resulting Wannier-Mott excitons [1, 2, 3, 4] have a radius that greatly exceeds the lattice constant $a$. Consequently, a continuum description suffices: only the states at the top of the valence band and at the bottom of the conduction band contribute significantly to the exciton wavefunction, and their dispersion is well approximated as quadratic. For three-dimensional (3D) crystals and for a $1/r$ Coulomb attraction, this maps the exciton spectrum to that of a hydrogen atom, with a rescaled Bohr radius and Rydberg energy [2, 3, 4, 5, 6].
@@
For a strongly screened short-range attraction with a characteristic scale $U$, excitons appear in 3D only if $U$ exceeds a critical value. However, in both two dimensions (2D) and one dimension (1D), a bound state exists for arbitrarily small $U$, with a binding energy $E_{\text{bind}}\propto\exp(-W/U)$ and $E_{\text{bind}}\propto U^{2}/W$, respectively [5, 8, 7] where $W$ is a measure of the average kinetic energy of the pair. These asymptotic estimates are widely used in standard discussions of excitons [9, 10] and provide a baseline for comparison between theory and experiments, for example, see Refs. 11, 12.
@@
However, many low-dimensional semiconductors that are technologically relevant host small, Frenkel and charge-transfer excitons [13, 14, 15] whose radii are equal or comparable to the lattice constant $a$. For example, TR-ARPES on epitaxial C60 films on Au(111) has revealed binding energies of 300–1000 meV for excitons that are either on the same C60 molecule or have the electron and hole on neighboring molecules [17, 16, 18, 19, 20]. These results agree with previous work on C60 using other optical techniques [21, 22]. Other relevant examples include organic crystals such as pentacene and PTCDA [23, 24], and magnetic insulators such as monolayer CrX3 [25, 26, 27, 28].
For such small, Frenkel-like excitons, the specific lattice symmetry becomes relevant beyond just the renormalization of the conduction and valence band effective masses [29]. This is because strong electron-hole attraction mixes hole (electron) states from across the entire valence (conduction) bands into the exciton wavefunction, so the details of the entire bands are now relevant. Furthermore, especially in organic semiconductors, the conduction and valence bands often have multi-orbital character, leading to spectral fingerprints and momentum-dependent symmetry changes [30, 31, 32].
This is why the study of small, Frenkel-like excitons requires (i) a lattice Hamiltonian that includes the relevant details of the valence and conduction bands, as well as a good description of the electron-hole attraction; and (ii) efficient methods to find the eigenenergies and eigenfunctions of eigenstates containing one particle-hole pair for this lattice Hamiltonian.
In this paper, we focus on aspect (ii), and propose a method adapted from studies of few-electron interacting systems [33] to calculate the real-space electron-hole propagators, from which the exciton spectrum and wavefunction can then be extracted. Although this method can be used for excitons of any size, its efficiency improves with decreasing exciton size, so it is most suitable for the study of small, Frenkel-like excitons. As further discussed below, we believe that in this limit, it offers a useful alternative to other methods for calculating exciton spectra, which are carried out in momentum space.
We exemplify this method by calculating exciton spectra for several model Hamiltonians in 1D and 2D, where the conduction and valence bands are described in terms of simple single- or multi-orbital tight-binding models, and the Coulomb attraction is assumed to be short-range. We note that these approximations are made for convenience, and we also discuss how they can be relaxed for more complex models. In terms of physics, we use these simple model Hamiltonians to investigate when and how the continuum approximation fails as the exciton's size approaches the lattice constant, and to highlight possible non-trivial behavior of small, Frenkel-like exciton spectra in systems with multi-orbital bands.
In principle, this method can also be used with realistic lattice Hamiltonians to calculate exciton spectra of specific materials, provided that accurate tight-binding models can be obtained from ab-initio methods and that the Coulomb attraction can be parameterized in a reasonably simple form, such as the Keldysh form [34, 35]. However, given the significant efforts and progress in developing efficient schemes for ab-initio simulations of excitons and their optical response, it is not clear to us whether our method would out-perform other approaches to solving the electron-hole Bethe-Salpeter equation, implemented in existing ab-initio codes [4, 35, 36, 37, 38, 39, 40]. The comparison might become favorable in the limit of small, Frenkel-like excitons, provided that the extraction of the parameters of the real-space Hamiltonian needed as the input for our method is not too computationally expensive.
In practice, we believe that the method we propose here is most suitable to study properties of small, Frenkel-like excitons in model Hamiltonians in order to gain intuition about the influence of the various terms of the Hamiltonian on their spectra and possible behaviors. These model Hamiltonians are an intermediary level of modeling between the simple continuum approximation (which works very well for Wannier-like excitons, but fails for small, Frenkel-like excitons) and full ab-initio simulations of excitons in specific materials.
Moreover, because our method is formulated in real space, one can generalize it beyond model Hamiltonians that are translationally invariant, to also study behavior of excitons near defects and/or interfaces which break translation symmetry. Another possible direction is to generalize the model Hamiltonian by adding Holstein or other types of model carrier-phonon couplings, to investigate their generic effects on the resulting exciton-polaron, as done in Ref. 41 for an extremely simple model.
@@
The article is organized as follows: Section II introduces the lattice model Hamiltonians we study, and Section III details our formalism. Our results are shown in Section IV, and we summarize our findings in Section V.
@@ II The lattice models
Here $n,m$ index lattice sites located at ${\bf R}_{n},{\bf R_{m}}$, $\alpha,\beta$ label the specific valence orbitals, and $v^{\dagger}_{\alpha n}$ creates an electron in the orbital $\alpha$ at site $n$. Explicitly, we will show results for 1D models with 2 orbitals per site ($1s$ and $2p_{x}$ for a chain parallel to the $x$ axis) and 2D triangular models with 3 orbitals per site ($d_{x^{2}-y^{2}},d_{xy},d_{3z^{2}-r^{2}}$ for a lattice lying in the $xy$ plane). Any other combinations of orbitals can be studied similarly. Crystal fields can be included trivially, but we ignore them for simplicity. We restrict the hopping to nearest-neighbor for simplicity, but this can be generalized straightforwardly to longer range. To minimize the number of free parameters for the 2D system, we use the Harrison rules [43] to calculate ratios between the hopping integrals, see Appendix A for details. For the same reasons, we present results only for a single-orbital conduction band:
@@
where $c_{n}^{\dagger}$ creates a conduction-band electron at site $n$ of the lattice. Here $\Delta$ is the difference between the on-site energies and controls the value of the band gap. For simplicity we consider again only nearest-neighbor hopping; generalization to longer range hopping is straightforward. Similarly, generalization to a multi-orbital conduction band is straightforward; this introduces more free parameters, but it can be dealt with similarly to how we treat the multi-orbital valence band.
We emphasize that here we ignore possible hybridization between the $c_{n}$ and $v_{\alpha m}$ orbitals, which can be a reasonable approximation for a large gap semiconductor under the right circumstances. Including such hybridization complicates matters because the real-space counterparts of the conduction and valence band Bloch eigenstates are no longer the atomic orbitals $c_{n}$ and $v_{\alpha m}$, but instead more extended Wannier orbitals. While dealing with such complications is beyond the scope of this work, we believe that they can be handled with the general method described below, and among other things, will give rise to an exchange interaction that will lift the degeneracy between singlet and triplet excitons which holds for the simpler model Hamiltonians discussed here. [40, 42] In the absence of such hybridization, the spin degrees of freedom are trivial, and from now on we drop the spin labels in this work.
@@
At first sight, this on-site attraction is an unreasonable approximation for organic semiconductors whose small, Frenkel-like excitons are precisely due to the very weakly screened long-range Coulomb interaction. Nevertheless, in the limit of on-site Frenkel excitons of most interest to us in this work, the electron and hole are on the same site with extremely high probability. Because they never 'explore' longer separations, the details of the longer-range Coulomb attraction become irrelevant. The only meaningful interaction energy scale in this limit is the difference between the attraction between an on-site pair and one occupying adjacent sites. This is what the $U_{\alpha}$ energies describe. For completeness, we mention that longer range Coulomb attraction can also be studied with the method we discuss below. A sample calculation for the effect of adding a nearest neighbor interaction can be found in Appendix C.
@@ III The method
Because of the translational invariance of the Hamiltonian (1), the exciton total momentum $\bf K$} is a good quantum number; its value is restricted to the Brillouin zone of the appropriate lattice. For any given value of $\bf K$}, a particularly convenient basis for calculating small, Frenkel-like exciton eigenstates is the set of all states [2, 9, 33]:
$|\alpha,{\bf K},\bm{\delta}\rangle=\sum_{n}{e^{i{\bf K}({\bf R}_{n}+\bm{\delta}/2)}\over\sqrt{N}}c^{\dagger}_{n}v_{\alpha,n+\bm{\delta}}|{\rm GS}\rangle$ (7)
where $|{\rm GS}\rangle$ is the ground state with filled valence bands and empty conduction band, and $N\rightarrow\infty$ is the number of lattice sites. This is because for small, Frenkel-like excitons, the relative displacement $\bm{\delta}$ between the electron and hole is short, meaning that the eigenfunction of an exciton with momentum $\bf K$}:
$|{\bf K},{\rm exc}\rangle=\sum_{\alpha,\bm{\delta}}\phi_{\bf K}(\alpha,\bm{\delta})|\alpha,{\bf K},\bm{\delta}\rangle$ (8)
will have significant amplitude $\phi_{\bf K}(\alpha,\bm{\delta})$ only for a few values of $\bm{\delta}$, like $\bm{\delta}=0$ when the electron and hole are at the same site, $|\bm{\delta}|=a$ where the electron and hole are on neighboring sites, etc. Because the electron and hole are bound, there is certainly a cutoff value $\delta_{M}$ beyond which the probability $|\phi_{\bf K}(\alpha,\bm{\delta})|^{2}$ to find the electron and hole at a displacement $|\bm{\delta}|\geq\delta_{M}$ becomes so small that it can be set to zero without loss of accuracy. The smaller the exciton, the smaller the value of $\delta_{M}$, and thus fewer coefficients $\phi_{\bf K}(\alpha,\bm{\delta})$ with $|\bm{\delta}|\leq\delta_{M}$ are needed to reconstruct an accurate exciton eigenfunction.
This should be contrasted with the more established approach of working entirely in the momentum space. In this case, the eigenfunction for an exciton with momentum $\bf K$} can be rewritten as:
$|{\bf K},{\rm exc}\rangle=\sum_{\alpha,{\bf k}}\phi_{\bf K}(\alpha,{\bf k}-{{\bf K}\over 2})c^{\dagger}_{\bf k}v_{\alpha,\bf{k}-\bf{K}}|{\rm GS}\rangle$ (9)
where $\phi_{\bf K}(\alpha,{\bf k}-{{\bf K}\over 2})=\sum_{\bm{\delta}}e^{i({\bf k}-{{\bf K}\over 2})\bm{\delta}}/\sqrt{N}\phi_{\bf K}(\alpha,\bm{\delta})$ is the Fourier transform of the real-space relative wavefunction $\phi_{\bf K}(\alpha,\bm{\delta})$. The more localized is the real space wavefunction $\phi_{\bf K}(\alpha,\bm{\delta})$, the more extended over the full Brillouin zone becomes $\phi_{\bf K}(\alpha,{\bf k}-{{\bf K}\over 2})$. In other words, to reconstruct an accurate small, Frenkel-like exciton eigenstate, one needs to calculate accurate values for $\phi_{\bf K}(\alpha,{\bf k}-{{\bf K}\over 2})$ in the entire Brillouin zone. Reaching convergence requires defining an appropriate grid of $\bf k$} points in the Brillouin zone; accuracy improves with an increasing number of grid-points, but this also increases computational costs. This is why for sufficiently small excitons, the real-space basis of Eq. 7 eventually becomes more efficient.
Conversely, as an exciton approaches the Wannier limit, working in the real-space basis becomes inefficient because $\phi_{\bf K}(\alpha,\bm{\delta})$ now spreads over many sites. In this limit, working in the momentum space becomes favorable because here $\phi_{\bf K}(\alpha,{\bf k}-{{\bf K}\over 2})$ is large only for momenta $\bf k$} near the bottom of conduction band.
We emphasize that the two approaches are fully equivalent and both will produce the same exciton spectra, the only difference being the computational cost to obtain accurate results. Next, we briefly explain how we find the exciton spectrum using the real-space basis.
Instead of solving the Schrödinger equation directly, it is more convenient to calculate instead the particle-hole real-space propagators [33]:
@@
$G_{\alpha,\beta}({\bf K},\bm{\delta},z)=\sum_{\lambda}\frac{\langle\alpha,{\bf K},\bm{0}|{\bf K},\lambda\rangle\langle{\bf K},\lambda|\beta,{\bf K},\bm{\delta}\rangle}{z-E({\bf K},\lambda)}$ (11)
@@
In the absence of Coulomb interactions, these electron-hole eigenstates are $c^{\dagger}_{\bf k}v_{{\bf k-K},\gamma}|{\rm GS}\rangle$ for any $\bf k$}, with the corresponding energy $E_{c}({\bf k})-E_{v}({\bf k-K},\gamma)$. They define the electron-hole continuum that for a given $\bf K$} starts at
$E_{\rm gap}({\bf K})=\min_{{\bf k},\gamma}[E_{c}({\bf k})-E_{v}({\bf k-K},\gamma)]$ (12)
@@
If the Coulomb attraction is sufficiently strong to bind excitons (which is always the case in 1D and 2D systems), their eigenenergies are signaled by discrete poles of $G_{\alpha,\beta}({\bf K},\bm{\delta},z)$ at energies inside the gap $\omega=E_{\rm exc}({\bf K})<E_{\rm gap}({\bf K})$. The presence of such discrete states signals stable bound electron-hole (excitons) eigenstates below the particle-hole continuum. The corresponding exciton binding energy is $E_{\rm bind}({\bf K})=E_{\rm gap}({\bf K})-E_{\rm exc}({\bf K})$ [4, 45].
@@
meaning that (see Eq. (8)):
$\phi^{*}_{\bf K}(\beta,\bm{\delta})=\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle\propto G_{\alpha,\beta}({\bf K},\bm{\delta},E_{\rm exc}({\bf K})).$ (14)
Eq. (14) gives the real-space exciton amplitude after normalization $\sum_{\beta,\bm{\delta}}|\phi_{\bf K}(\beta,\bm{\delta})|^{2}=1$. We also use it to extract the exciton radius $\xi$ from the long-range envelope $\phi_{\bf K}(\beta,\bm{\delta})\rightarrow\exp(-|\bm{\delta}|/\xi)$ when $|\bm{\delta}|\rightarrow\infty$ [46].
These real-space particle-hole propagators are calculated from matrix elements between states of Eq.  (7) of the identity $\hat{G}(z)(z-{\cal H})=1$, from which we find:
@@
The second term can be expanded in terms of other real-space propagators, in particular for models with only nearest-neighbor hopping considered here, the right hand side will also contain propagators where the electron-hole displacement is either increased or decreased by one lattice constant, due to hopping. We can then generate a linear system of coupled equations linking all these propagators, from which we can calculate their values for any energy. In principle, the size of this system is infinite because $\bm{\delta}$ can be arbitrarily large, but as already discussed, for small, Frenkel-like excitons the magnitude of the propagators corresponding to energies inside the gap decreases very fast with increasing $\bm{\delta}$ [46]. This allows us to truncate this system by setting all propagators to zero with $|\bm{\delta}|\geq\delta_{M}$, where the cutoff $\delta_{M}$ is increased until the exciton energy is converges within the desired accuracy. For very small, Frenkel excitons, as shown below, reasonable accuracy may be obtained even for a cutoff $\delta_{M}=2a$, making such calculations extremely efficient. More details and the explicit equations for these propagators are given in Appendix B.
Figure 1: Binding energy of the 1D exciton with momentum K=0{K}=0 (left panel) and its radius in units $a=1$ (right panel) from the 1D lattice model (red circles) with a one-orbital conduction band with $t_{c}=1$, and a $s,p_{x}$ two-orbital valence band with $t_{ss}=1.1,t_{pp}=-1,|t_{sp}|=0.5$. The corresponding valence band structure is shown in Fig. 2. Here, we assume $U_{s}=U_{p}=U$. The blue line shows the continuum approximation predictions. The vertical dashed line marks the value of $U$ above which the continuum approximation starts to fail. See text for more details.